# Jiaming Wei    Dihan Zou

### v2

@@ 1 Introduction
Classical models of monopoly regulation assume that regulators can freely deploy the full range of policy instruments including taxes, subsidies, and price caps. In practice, however, these instruments are often restricted. Such restrictions complicate the design of optimal regulation, especially when the regulator lacks full information about production costs. In many environments, subsidies are infeasible due to institutional rules, political externalities, or fiscal constraints. In the European Union, for instance, state aid to firms is tightly restricted, with violations triggering substantial penalties, as in the case FrancevEC. [^1] France enticed the dominant operator of ferry travel to and from Corsica to offer a richer schedule with a per capita subsidy, but the European Commission considered it inadmissible. The operator, Société Nationale Corse-Méditerranée, had to repay €220 million of illegal state aid (legalNotes). Subsidizing goods that enter export markets can provoke trade disputes, [^2] See EUTradeWar. In principle, WTOsubsidy prohibits subsidizing exports. while in developing countries limited public budgets severely constrain governments’ scope of subsidy. [^3] In many developing countries, governments lack sufficient public funds to subsidize utilities like electricity (UNReportSubsidy). These restrictions raise a natural question: How should a monopolist with private cost information be regulated when subsidies are prohibited?
This paper studies monopoly regulation under asymmetric cost information when subsidies are infeasible. We adopt the seminal framework of BaronMyerson82: a monopolist serves a single-product market and posts a price, facing downward-sloping demand. The firm’s marginal cost is privately observed. The regulator maximizes a weighted welfare function combining consumer surplus and firm profit. Crucially, the regulator cannot use subsidies. Without loss of generality, we focus on unit taxes as the sole policy instrument, which are equivalent to lump-sum excise taxes but provide a natural economic interpretation. A unit tax increases the consumer price relative to the firm’s price, reducing demand. By designing a tax schedule that maps prices into tax rates, the regulator can influence pricing incentives. The policy faces a central tradeoff: absent intervention, the monopolist sets its unregulated monopoly price, generating inefficiently low output; yet while taxes can deter excessive firm prices, they also raise consumer prices and thus depress demand and gains from trade.
@@
When intervention is necessary, we find simple sufficient conditions for the optimal policy to be a progressive price cap. Under this policy, no tax applies to prices below a benchmark level, while prices above this threshold incur a positive tax that is increasing in price, which can be prohibitive for sufficiently high prices. The benchmark price partitions the price space into two regions: a delegation region, where the monopolist retains pricing power, and a taxation region, where high prices are deterred. The firm self-selects across these regions based on its cost: a low-cost firm either chooses laissez-faire prices or the benchmark price, a medium-cost firm pools at the benchmark, and a high-cost firm charges above it, facing taxes and, in extreme cases, exclusion. This structure echoes the carrot-and-stick logic of BaronMyerson82, but without subsidies the regulator cannot reward low prices with transfers; instead, she offers the weaker reward of price-setting freedom. The design of the progressive price cap reflects a three-way tradeoff among access, affordability, and profitability, which are distinct forces in the regulator’s objective. Access concerns how many consumers are ultimately served, capturing the extensive margin of consumer surplus. Affordability refers to the prices paid by those who do purchase, i.e., the intensive margin. Profitability reflects the firm’s ability to cover its costs and remain willing to produce. Improving affordability requires taxing high prices, but doing so makes high-cost firms less profitable and reduces access, since the resulting consumer price moves farther from efficiency. The benchmark price and the accompanying tax schedule in the progressive price cap are chosen to balance these competing objectives.
@@
This paper contributes to the vast literature on monopoly regulation with asymmetric information. We refer to LaffontTiroleBook93 and ArmstrongSappington07HIO for an overview of the field. Classical literature begins by modeling different modes of regulator’s information disadvantage, including unknown cost (BaronMyerson82), unknown demand (LewisSappington88AER), and hidden cost-reduction effort (LaffontTirole86JPE). Subsequent literature has been mainly focusing on more severe asymmetric information (LewisSappington88unknown-demand-and-cost; Armstrong99JET) toward larger uncertainty and culminates in robustness (GuoShmaya25RobustRegu; Garrett14GEB; MishraPatil2025JET; MishraPatilPavan) and flexibility (Krahmer25Flexible). The mode of asymmetric information in our setting is closest to BaronMyerson82. The regulator is uncertain about the firm’s marginal cost, but has a prior distribution for policy design. We extend the literature by introducing policy restrictions alongside informational asymmetry, an aspect that has received comparatively less attention. [^4] There is a literature evaluating different policy instruments under specific applications such as pollution abatement (RobertsSpence1976, Baron85RAND, Baron85JPubE). A recent contribution by BFV17RANDPriceQuantity compares price control and quantity control under different modes of information asymmetry. Pioneered by Schmalensee89Good, the literature on policy restriction focuses on the extreme case of no transfer and highlights the performance of simple instruments (e.g., BhaskarMcClellanAER23; Xia25).
Closest to our work is AmadorBagwell22Regulation, who establish optimality of price cap in general market environments under asymmetric information about production cost without transfers, extending the seminal Lagrangian technique in AB13ECMA. We bridge the extremes of flexible transfers and no transfer by analyzing an intermediate case: subsidies are prohibited, but taxes remain feasible. To our best knowledge, this is the first paper to study optimal monopoly regulation under asymmetric information with one-sided transfers. [^5] In Section 2.7 of their textbook (pp. 151 - 153), LaffontTiroleBook93 discuss a setting of hidden cost-reduction effort with a budget balance constraint for the firm, and suggest the optimum may be implemented using a “sliding scale plan.”
@@ 2 Environment
In a single-product market, a monopolistic firm produces quantity $q\in[0,q_{max}]$ at a constant marginal cost $c$. For $q>0$, it incurs a fixed cost $k\geq 0$, so its cost function is $C(q)=c\cdot q+k$; otherwise $C(0)=0$. The marginal cost $c$ is drawn from a twice continuously differentiable distribution $F$ on $[0,1]$, and the realization of $c$, henceforth the firm’s type, is privately observed by the firm. The fixed cost $k$ is commonly known, ensuring that asymmetric information pertains solely to the marginal cost relevant for the firm’s output decision. [^6] This specification also maintains comparability with the canonical models in BaronMyerson82 and AmadorBagwell22Regulation. The firm posts a unit price $p$.
@@
One type of regulation commonly seen in the literature (e.g., BaronMyerson82) manipulates the demand curve so that it is more sensitive (elastic) to price. The following example illustrates how different choices of unit tax shape the demand curve.
@@
$=\underbrace{\int_{0}^{q_{\tau}(p)}P(x)\,dx-cq_{\tau}(p)}_{\text{trading surplus}}\quad-\quad(1-\alpha)\cdot\Pi_{\tau}(c).$
@@
$\max_{q_{\tau}\colon\mathbb{R_{+}}\rightarrow[0,q_{max}]}\int_{0}^{1}\left[V(q_{\tau}(p(c)))-cq_{\tau}(p(c))-(1-\alpha)\cdot\Pi_{\tau}(c)\right]dF(c),$ (2)
@@ 2.1 Preliminary analysis
$\max_{p,q}\int_{0}^{1}\left(V(q(c))-cq(c)-(1-\alpha)\Pi(c)\right)dF(c)$ (5)
@@
This new dichotomy between monetary and allocative distortions makes the optimal regulation markedly different from the classical results. Traditional monopoly regulation with flexible transfers (e.g., BaronMyerson82) incentivizes low prices by granting additional demand (with subsidies) to low-cost reports and reducing demand (with taxes) for high-cost reports. The resulting regulation is distortionary in demand relative to ex post efficiency (marginal cost pricing: $P(q(c))=c$).
@@ 2.2 Discussion
We restrict the regulator to using a unit tax, ruling out piece-rate or lump-sum subsidies. As shown in Lemma B.1, an excise (lump-sum) tax is equivalent to a unit tax in our environment. Without the transfer constraint, the regulator could employ a fully flexible policy combining both taxes and subsidies to manipulate the price sensitivity of the demand curve and intervene at virtually all price levels. Our formulation is therefore equivalent to BaronMyerson82 with an additional no-subsidy constraint, and also equivalent to AmadorBagwell22Regulation once their “no-tax” restriction is removed (see Lemma B.1).
@@
Finally, we assume that the revenue function induced by market demand is strictly concave, that the demand curve reaches both its upper and lower bounds, and that the distribution of cost types is twice continuously differentiable. These assumptions are stronger than those in BaronMyerson82 and weaker than those in AmadorBagwell22Regulation. We regard them as mild: they still accommodate a wide range of market primitives while ensuring a well-behaved laissez-faire pricing schedule. They also deliver continuity, existence, and uniqueness of the optimal regulatory policy. In Lemma B.4, we show that it is without loss of generality to restrict attention to deterministic policies.
@@ 3.1 Laissez-faire as the optimal regulation
When $\alpha\in[0,1)$, the regulator places extra weight on consumer surplus. Condition (LF) can then be interpreted as a weighted monopoly markup adjusted by a redistributive term, $(1-\alpha)F(c)$. As the regulator becomes more consumer-oriented (i.e., as $\alpha$ decreases), the desire to reduce prices strengthens, and it becomes harder to justify laissez-faire. [^10] In footnote 32, AmadorBagwell22Regulation (p. 1743) note that their exclusion threshold weakly increases in $\alpha$.
@@
Condition (LF) is complementary to the sufficient conditions for nonintervention in AmadorBagwell22Regulation, Proposition 3.111111Their sufficient conditions are: (a’) $qP^{\prime}(q)$ is decreasing in $q$, and (b) $f(c)$ is nondecreasing. Condition (a’) implies the monopoly markup $q_{LF}(c)P^{\prime}(q_{LF}(c))$ decreases in $q_{LF}(c)$, equivalent to part (a) in (LF) increasing in $c$. When $\alpha=1$, their conditions imply ours. For general $\alpha\in[0,1)$, the two sets of conditions are complementary. The regulatory environment they study—monopoly regulation without transfers—is equivalent to ours with an additional no-tax constraint. There, the regulator can use only a price cap to exclude high-cost types. This instrument can be replicated through taxation in our framework (cf. policy example $\tau_{1}$). Naturally, if it is optimal not to intervene via taxes (i.e., without subsidies) in a given market environment, then it is also optimal not to intervene via a price cap (i.e., without transfers). Thus, (LF) serves as a sufficient condition for laissez-faire to be optimal when transfers are prohibited. Importantly, allowing taxes expands the regulator’s intervention set relative to the no-transfer case. The underlying tradeoff is similar, though: reducing prices versus distorting high-cost types.
At the opposite extreme, when the regulator can use fully flexible transfers, BaronMyerson82 show that intervention is typically desirable. From a feasibility perspective, laissez-faire is always feasible regardless of transfer restrictions. With flexible transfers, however, laissez-faire is sub-optimal except for knife-edge cases. Introducing the no-subsidy constraint reduces the feasible set in a way that can make nonintervention optimal, as shown in Proposition 1. From the perspective of policy instruments, Proposition 1 also shows that prohibiting subsidies limits the regulator’s ability to use taxes effectively.
@@
The preceding discussion suggests that taxes and subsidies function both as substitute and as complementary instruments. They are substitute in that either can be used to induce lower prices, albeit at the cost of distortion. But they are complementary because the distortions they induce move in different directions: taxes reduce output, while subsidies entail fiscal costs (when $\alpha<1$). When both tools are available, the regulator balances these distortions—using taxes and subsidies jointly to maintain low prices without excessively depressing output or incurring excessive expenditure. This complementarity reduces overall welfare loss. In contrast, when transfers are infeasible, the absence of the complementary instruments often renders the welfare loss too large to justify intervention (AmadorBagwell22Regulation, Proposition 3).
@@
Note that $A-c]^{+}$ is a nonnegative decreasing function in $c$. If $f(c)$ is nonincreasing (e.g., uniform and truncated exponential), the whole expression is nonincreasing. It is also easy to check that the expression strictly decreases in some intervals for truncated normal densities. We conclude that there is room of intervention under linear demand curve and many common cost distributions, regardless the welfare weight $\alpha$. In contrast, without any transfer, laissez-faire will be optimal under the environment of linear demand, uniform cost distribution, and $\alpha=1$. [^13] See AmadorBagwell22Regulation’s result without fixed cost (their footnote 33, p. 1744). For a concrete example on the substitution relationship between taxes and subsidies, we compare the optimal policy in Section 3.2 for this particular environment across three policy sets: only taxation, no transfer, and flexible transfer. [^14] We henceforth refer to BaronMyerson82 whenever discussing on the optimal policies with flexible transfer, and refer to AmadorBagwell22Regulation whenever discussing on the optimal policies with no transfer.

### v3

@@ 1 Introduction
Classical models of monopoly regulation assume that regulators can freely deploy the full range of policy instruments including taxes, subsidies, and price caps. In practice, however, these instruments are often restricted. Such restrictions complicate the design of optimal regulation, especially when the regulator lacks full information about production costs. In many environments, subsidies are infeasible due to institutional rules, political externalities, or fiscal constraints. In the European Union, for instance, state aid to firms is tightly restricted, with violations triggering substantial penalties, as in the case French Republic v European Commision (2017). [^1] France enticed the dominant operator of ferry travel to and from Corsica to offer a richer schedule with a per capita subsidy, but the European Commission considered it inadmissible. The operator, Société Nationale Corse-Méditerranée, had to repay €220 million of illegal state aid (Giraud (2017)). Subsidizing goods that enter export markets can provoke trade disputes, [^2] See European Commission (2013). In principle, WTO (1994) prohibits subsidizing exports. while in developing countries limited public budgets severely constrain governments’ scope of subsidy. [^3] In many developing countries, governments lack sufficient public funds to subsidize utilities like electricity (Kim and Horn (1999)). These restrictions raise a natural question: How should a monopolist with private cost information be regulated when subsidies are prohibited?
This paper studies monopoly regulation under asymmetric cost information when subsidies are infeasible. We adopt the seminal framework of Baron and Myerson (1982): a monopolist serves a single-product market and posts a price, facing downward-sloping demand. The firm’s marginal cost is privately observed. The regulator maximizes a weighted welfare function combining consumer surplus and firm profit. Crucially, the regulator cannot use subsidies. Without loss of generality, we focus on unit taxes as the sole policy instrument, which are equivalent to lump-sum excise taxes but provide a natural economic interpretation. A unit tax increases the consumer price relative to the firm’s price, reducing demand. By designing a tax schedule that maps prices into tax rates, the regulator can influence pricing incentives. The policy faces a central tradeoff: absent intervention, the monopolist sets its unregulated monopoly price, generating inefficiently low output; yet while taxes can deter excessive firm prices, they also raise consumer prices and thus depress demand and gains from trade.
@@
When intervention is necessary, we find simple sufficient conditions for the optimal policy to be a progressive price cap. Under this policy, no tax applies to prices below a benchmark level, while prices above this threshold incur a positive tax that is increasing in price, which can be prohibitive for sufficiently high prices. The benchmark price partitions the price space into two regions: a delegation region, where the monopolist retains pricing power, and a taxation region, where high prices are deterred. The firm self-selects across these regions based on its cost: a low-cost firm either chooses laissez-faire prices or the benchmark price, a medium-cost firm pools at the benchmark, and a high-cost firm charges above it, facing taxes and, in extreme cases, exclusion. This structure echoes the carrot-and-stick logic of Baron and Myerson (1982), but without subsidies the regulator cannot reward low prices with transfers; instead, she offers the weaker reward of price-setting freedom. The design of the progressive price cap reflects a three-way tradeoff among access, affordability, and profitability, which are distinct forces in the regulator’s objective. Access concerns how many consumers are ultimately served, capturing the extensive margin of consumer surplus. Affordability refers to the prices paid by those who do purchase, i.e., the intensive margin. Profitability reflects the firm’s ability to cover its costs and remain willing to produce. Improving affordability requires taxing high prices, but doing so makes high-cost firms less profitable and reduces access, since the resulting consumer price moves farther from efficiency. The benchmark price and the accompanying tax schedule in the progressive price cap are chosen to balance these competing objectives.
@@
This paper contributes to the vast literature on monopoly regulation with asymmetric information. We refer to Laffont and Tirole (1993) and Armstrong and Sappington (2007) for an overview of the field. Classical literature begins by modeling different modes of regulator’s information disadvantage, including unknown cost (Baron and Myerson (1982)), unknown demand (Lewis and Sappington (1988b)), and hidden cost-reduction effort (Laffont and Tirole (1986)). Subsequent literature has been mainly focusing on more severe asymmetric information (Lewis and Sappington (1988a); Armstrong (1999)) toward larger uncertainty and culminates in robustness (Guo and Shmaya (2025); Garrett (2014); Mishra and Patil (2025); Mishra et al. (2025)) and flexibility (Krähmer (2025)). The mode of asymmetric information in our setting is closest to Baron and Myerson (1982). The regulator is uncertain about the firm’s marginal cost, but has a prior distribution for policy design. We extend the literature by introducing policy restrictions alongside informational asymmetry, an aspect that has received comparatively less attention. [^4] There is a literature evaluating different policy instruments under specific applications such as pollution abatement (Roberts and Spence (1976), Baron (1985a), Baron (1985b)). A recent contribution by Basso et al. (2017) compares price control and quantity control under different modes of information asymmetry. Pioneered by Schmalensee (1989), the literature on policy restriction focuses on the extreme case of no transfer and highlights the performance of simple instruments (e.g., Bhaskar et al. (2023); Xia (2025)).
Closest to our work is Amador and Bagwell (2022), who establish optimality of price cap in general market environments under asymmetric information about production cost without transfers, extending the seminal Lagrangian technique in Amador and Bagwell (2013). We bridge the extremes of flexible transfers and no transfer by analyzing an intermediate case: subsidies are prohibited, but taxes remain feasible. To our best knowledge, this is the first paper to study optimal monopoly regulation under asymmetric information with one-sided transfers. [^5] In Section 2.7 of their textbook (pp. 151 - 153), Laffont and Tirole (1993) discuss a setting of hidden cost-reduction effort with a budget balance constraint for the firm, and suggest the optimum may be implemented using a “sliding scale plan.”
@@ 2 Environment
In a single-product market, a monopolistic firm produces quantity $q\in[0,q_{max}]$ at a constant marginal cost $c$. For $q>0$, it incurs a fixed cost $k\geq 0$, so its cost function is $C(q)=c\cdot q+k$; otherwise $C(0)=0$. The marginal cost $c$ is drawn from a twice continuously differentiable distribution $F$ on $[0,1]$, and the realization of $c$, henceforth the firm’s type, is privately observed by the firm. The fixed cost $k$ is commonly known, ensuring that asymmetric information pertains solely to the marginal cost relevant for the firm’s output decision. [^6] This specification also maintains comparability with the canonical models in Baron and Myerson (1982) and Amador and Bagwell (2022). The firm posts a unit price $p$.
@@
One type of regulation commonly seen in the literature (e.g., Baron and Myerson (1982)) manipulates the demand curve so that it is more sensitive (elastic) to price. The following example illustrates how different choices of unit tax shape the demand curve.
@@
$=\underbrace{\int_{0}^{q_{\tau}(p)}P(x)\,dx-cq_{\tau}(p)-k\cdot\mathbbm{1}_{q_{\tau}(p)>0}}_{\text{trading surplus}}\quad-\quad(1-\alpha)\cdot\Pi_{\tau}(c).$
@@
$\max_{q_{\tau}\colon\mathbb{R_{+}}\rightarrow[0,q_{max}]}\int_{0}^{1}\left[V(q_{\tau}(p(c)))-cq_{\tau}(p(c))-k\cdot\mathbbm{1}_{q_{\tau}(p)>0}-(1-\alpha)\cdot\Pi_{\tau}(c)\right]dF(c),$ (2)
@@ 2.1 Preliminary analysis
$\max_{p,q}\int_{0}^{1}\left(V(q(c))-cq(c)-k\cdot\mathbbm{1}_{q(c)>0}-(1-\alpha)\Pi(c)\right)dF(c)$ (5)
@@
This new dichotomy between monetary and allocative distortions makes the optimal regulation markedly different from the classical results. Traditional monopoly regulation with flexible transfers (e.g., Baron and Myerson (1982)) incentivizes low prices by granting additional demand (with subsidies) to low-cost reports and reducing demand (with taxes) for high-cost reports. The resulting regulation is distortionary in demand relative to ex post efficiency (marginal cost pricing: $P(q(c))=c$).
@@ 2.2 Discussion
We restrict the regulator to using a unit tax, ruling out piece-rate or lump-sum subsidies. As shown in Lemma B.1, an excise (lump-sum) tax is equivalent to a unit tax in our environment. Without the transfer constraint, the regulator could employ a fully flexible policy combining both taxes and subsidies to manipulate the price sensitivity of the demand curve and intervene at virtually all price levels. Our formulation is therefore equivalent to Baron and Myerson (1982) with an additional no-subsidy constraint, and also equivalent to Amador and Bagwell (2022) once their “no-tax” restriction is removed (see Lemma B.1).
@@
Finally, we assume that the revenue function induced by market demand is strictly concave, that the demand curve reaches both its upper and lower bounds, and that the distribution of cost types is twice continuously differentiable. These assumptions are stronger than those in Baron and Myerson (1982) and weaker than those in Amador and Bagwell (2022). We regard them as mild: they still accommodate a wide range of market primitives while ensuring a well-behaved laissez-faire pricing schedule. They also deliver continuity, existence, and uniqueness of the optimal regulatory policy. In Lemma B.4, we show that it is without loss of generality to restrict attention to deterministic policies.
@@ 3.1 Laissez-faire as the optimal regulation
When $\alpha\in[0,1)$, the regulator places extra weight on consumer surplus. Condition (LF) can then be interpreted as a weighted monopoly markup adjusted by a redistributive term, $(1-\alpha)F(c)$. As the regulator becomes more consumer-oriented (i.e., as $\alpha$ decreases), the desire to reduce prices strengthens, and it becomes harder to justify laissez-faire. [^10] In footnote 32, Amador and Bagwell (2022) (p. 1743) note that their exclusion threshold weakly increases in $\alpha$.
@@
Condition (LF) is complementary to the sufficient conditions for nonintervention in Amador and Bagwell (2022), Proposition 3.111111Their sufficient conditions are: (a’) $qP^{\prime}(q)$ is decreasing in $q$, and (b) $f(c)$ is nondecreasing. Condition (a’) implies the monopoly markup $q_{LF}(c)P^{\prime}(q_{LF}(c))$ decreases in $q_{LF}(c)$, equivalent to part (a) in (LF) increasing in $c$. When $\alpha=1$, their conditions imply ours. For general $\alpha\in[0,1)$, the two sets of conditions are complementary. The regulatory environment they study—monopoly regulation without transfers—is equivalent to ours with an additional no-tax constraint. There, the regulator can use only a price cap to exclude high-cost types. This instrument can be replicated through taxation in our framework (cf. policy example $\tau_{1}$). Naturally, if it is optimal not to intervene via taxes (i.e., without subsidies) in a given market environment, then it is also optimal not to intervene via a price cap (i.e., without transfers). Thus, (LF) serves as a sufficient condition for laissez-faire to be optimal when transfers are prohibited. Importantly, allowing taxes expands the regulator’s intervention set relative to the no-transfer case. The underlying tradeoff is similar, though: reducing prices versus distorting high-cost types.
At the opposite extreme, when the regulator can use fully flexible transfers, Baron and Myerson (1982) show that intervention is typically desirable. From a feasibility perspective, laissez-faire is always feasible regardless of transfer restrictions. With flexible transfers, however, laissez-faire is sub-optimal except for knife-edge cases. Introducing the no-subsidy constraint reduces the feasible set in a way that can make nonintervention optimal, as shown in Proposition 1. From the perspective of policy instruments, Proposition 1 also shows that prohibiting subsidies limits the regulator’s ability to use taxes effectively.
@@
The preceding discussion suggests that taxes and subsidies function both as substitute and as complementary instruments. They are substitute in that either can be used to induce lower prices, albeit at the cost of distortion. But they are complementary because the distortions they induce move in different directions: taxes reduce output, while subsidies entail fiscal costs (when $\alpha<1$). When both tools are available, the regulator balances these distortions—using taxes and subsidies jointly to maintain low prices without excessively depressing output or incurring excessive expenditure. This complementarity reduces overall welfare loss. In contrast, when transfers are infeasible, the absence of the complementary instruments often renders the welfare loss too large to justify intervention (Amador and Bagwell (2022), Proposition 3).
@@
Note that $A-c]^{+}$ is a nonnegative decreasing function in $c$. If $f(c)$ is nonincreasing (e.g., uniform and truncated exponential), the whole expression is nonincreasing. It is also easy to check that the expression strictly decreases in some intervals for truncated normal densities. We conclude that there is room of intervention under linear demand curve and many common cost distributions, regardless the welfare weight $\alpha$. In contrast, without any transfer, laissez-faire will be optimal under the environment of linear demand, uniform cost distribution, and $\alpha=1$. [^13] See Amador and Bagwell (2022)’s result without fixed cost (their footnote 33, p. 1744). For a concrete example on the substitution relationship between taxes and subsidies, we compare the optimal policy in Section 3.2 for this particular environment across three policy sets: only taxation, no transfer, and flexible transfer. [^14] We henceforth refer to Baron and Myerson (1982) whenever discussing on the optimal policies with flexible transfer, and refer to Amador and Bagwell (2022) whenever discussing on the optimal policies with no transfer.